OpenStudy (anonymous):

(5y^2)/(3dk^2)+(ku^3)/(6a^2d^2) simplify as much as possible.

OpenStudy (mertsj):

\[\frac{5y^2}{3dk^2}+\frac{ku^3}{6a^2d^2}=\frac{10a^2y^2+k^3u^3}{6a^2d^2k^2}\]